The personnel office at a large electronics firm regularly schedules job interviews and maintains records of the interviews. From the past records, they have found that the length of a first interview is normally distributed, with mean μ = 36 minutes and standard deviation σ = 7 minutes. (Round your answers to four decimal places.)
Answer #1

mean = \mu = 36

standard deviation = \sigma = 7

The probability that a first interview will last 40 minutes or longer is,

P(x \geq 40) = 1 - P(x  \leq 40)

= 1 - P[(x - \mu ) / \sigma \leq (40 - 36) / 7]

= 1 -  P(z \leq 0.57)  

= 1 - 0.7157

= 0.2843